# Differentially Private Two-Stage Gradient Descent for Instrumental Variable Regression

### v1

@@ 4.1 Synthetic Data Simulations
We generate synthetic data according to the IVaR model in (1). To simulate the correlation between $\bm{\epsilon}_{1}$ and $\epsilon_{2}$, we include a counfounder $\mathbf{u}\in\mathbb{R}^{r}$, and set $\epsilon_{1}=\bm{\Phi}^{\top}\mathbf{u}_{i}+\bm{\epsilon}_{x}$ and $\epsilon_{2}=\bm{\phi}^{\top}\mathbf{u}+\epsilon_{y}$, and generate each entry of the dataset $(\mathbf{Z},\mathbf{X},\mathbf{Y})=\{(\mathbf{z}_{i},\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ according to the following model: $\mathbf{x}_{i}=\bm{\Theta}^{\top}\mathbf{z}_{i}+\bm{\Phi}^{\top}\mathbf{u}_{i}+\bm{\epsilon}_{x,i},$ and $\mathbf{y}_{i}=\bm{\beta}^{\top}\mathbf{x}_{i}+\bm{\phi}^{\top}\mathbf{u}_{i}+\epsilon_{y,i},$ where the ground-truth parameters are $\bm{\beta}\in\mathbb{R}^{p},\bm{\Theta}\in\mathbb{R}^{q\times p}$, $\mathbf{\Phi}\in\mathbb{R}^{r\times p}$, $\bm{\phi}\in\mathbb{R}^{r}$. These parameters are drawn as follows: $\bm{\beta}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{p})$, $\bm{\Theta}\sim 5\mathbf{I}_{q\times p}+\mathbf{E}$ with $\mathbf{E}_{ij}\sim\mathcal{N}(0,1)$. $\bm{\Phi}_{ij}\sim\mathcal{N}(0,1)$, and $\bm{\phi}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{r})$. For each simulation, we then sample $\mathbf{z}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{q})$, $\mathbf{u}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{r})$, $\bm{\epsilon}_{x,i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{p})$, and $\epsilon_{y,i}\sim\mathcal{N}(0,1)$.
@@ 4.2 Real-Data Experiments
We further evaluate our algorithm on the Angrist dataset (Angrist and Evans, 1998), which has been widely applied in the IVaR literature. This study examines the causal effect of fertility on female labor supply, leveraging the gender composition of the first two children as an instrument [^2] Research shows that parents whose first two children are of the same sex are significantly more likely to have an additional child (Westoff and Parke, 1972). At the same time, the sex composition of the first two children can be treated as randomly assigned and is not directly related to the mother’s labor supply.. The endogenous regressor $\mathbf{x}$ is the number of children bearing, the outcome $\mathbf{y}$ is the mother’s labor supply measured in number of working weeks per year, and the instrument $\mathbf{z}$ is a binary variable indicating whether the first two children are of the same gender. The original dataset contains $394,835$ samples. For illustration purpose, we randomly draw a subset of $20,000$ samples and keep $n=8065$ effective observations with number of children $\geq 2$. We center all variables $\mathbf{z},\mathbf{x},\mathbf{y}$ and run Algorithm 1 with $T=20$ iterations. Figure 5 presents the results over 1000 independent runs with privacy budgets $\rho_{1}=10,\rho_{2}=10$. As shown in Figure 5(a), the estimated $\bm{\beta}^{(T)}$ concentrates around 4.3-4.3, indicating that having an additional child reduces the mother’s labor supply by approximately 4.3 weeks per year. This estimate is consistent with the 2SLS benchmark.
@@ Appendix G Additional Experiments
G.1 Convergence Rate Comparison
@@
G.2 Additional Experiments on Angrist Dataset
We provide additional experimental results on the Angrist dataset with different privacy parameters $\rho_{1},\rho_{2}$. We consider two settings of privacy parameters: (i) $\rho_{1}=1,\rho_{2}=1$; (ii) $\rho_{1}=100,\rho_{2}=100$. The results are shown in Figures 8 and 9. We observe that when $\rho_{1},\rho_{2}$ are small, the estimates of $\bm{\beta}^{(T)}$ have larger variance. When $\rho_{1},\rho_{2}$ are larger, the estimates of $\bm{\beta}^{(T)}$ are more concentrated around the expected value. In both settings, the estimates of $\bm{\beta}^{(t)}$ converge in expectation within $T=20$ iterations.
Figure 9: Results on the Angrist dataset with $T=20,\rho_{1}=100,\rho_{2}=100$. (a) Boxplot of estimated $\bm{\beta}^{(T)}$, over 1000 runs. (b) Learning paths of parameters $\bm{\beta}^{(t)},\bm{\Theta}^{(t)}$, over 1000 runs. The shaded area represents the standard error.

### v2

@@ 4.1 Synthetic Data Simulations
We generate synthetic data according to the IVaR model in (1). To simulate the correlation between $\bm{\epsilon}_{1}$ and $\epsilon_{2}$, we include a confounder $\mathbf{u}\in\mathbb{R}^{r}$, and set $\epsilon_{1}=\bm{\Phi}^{\top}\mathbf{u}_{i}+\bm{\epsilon}_{x}$ and $\epsilon_{2}=\bm{\phi}^{\top}\mathbf{u}+\epsilon_{y}$, and generate each entry of the dataset $(\mathbf{Z},\mathbf{X},\mathbf{Y})=\{(\mathbf{z}_{i},\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ according to the following model: $\mathbf{x}_{i}=\bm{\Theta}^{\top}\mathbf{z}_{i}+\bm{\Phi}^{\top}\mathbf{u}_{i}+\bm{\epsilon}_{x,i},$ and $\mathbf{y}_{i}=\bm{\beta}^{\top}\mathbf{x}_{i}+\bm{\phi}^{\top}\mathbf{u}_{i}+\epsilon_{y,i},$ where the ground-truth parameters are $\bm{\beta}\in\mathbb{R}^{p},\bm{\Theta}\in\mathbb{R}^{q\times p}$, $\mathbf{\Phi}\in\mathbb{R}^{r\times p}$, $\bm{\phi}\in\mathbb{R}^{r}$. These parameters are drawn as follows: $\bm{\beta}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{p})$, $\bm{\Theta}\sim 5\mathbf{I}_{q\times p}+\mathbf{E}$ with $\mathbf{E}_{ij}\sim\mathcal{N}(0,1)$. $\bm{\Phi}_{ij}\sim\mathcal{N}(0,1)$, and $\bm{\phi}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{r})$. For each simulation, we then sample $\mathbf{z}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{q})$, $\mathbf{u}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{r})$, $\bm{\epsilon}_{x,i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{p})$, and $\epsilon_{y,i}\sim\mathcal{N}(0,1)$.
@@ 4.2 Real-Data Experiments
We further evaluate our algorithm on the Angrist dataset (Angrist and Evans, 1998), which has been widely applied in the IVaR literature. This study examines the causal effect of children bearing on female labor supply, leveraging the gender composition of the first two children as an instrument [^4] Research shows that parents whose first two children are of the same sex are significantly more likely to have an additional child (Westoff and Parke, 1972). At the same time, the sex composition of the first two children can be treated as randomly assigned and is not directly related to the mother’s labor supply.. The endogenous regressor $\mathbf{x}$ is the number of children bearing, the outcome $\mathbf{y}$ is the mother’s labor supply measured in number of working weeks per year, and the instrument $\mathbf{z}$ is a binary variable indicating whether the first two children are of the same gender. The original dataset contains $394,835$ samples. For illustration purpose, we randomly draw a subset of $20,000$ samples and keep $n=8065$ effective observations with number of children $\geq 2$. We center all variables $\mathbf{z},\mathbf{x},\mathbf{y}$ and run Algorithm 1 with $T=20$ iterations. Figure 5 presents the results over 1000 independent runs with privacy budgets $\rho_{1}=1,\rho_{2}=1$. As shown in Figure 5(a), the estimated $\bm{\beta}^{(T)}$ concentrates around 4.3-4.3, indicating that having an additional child reduces the mother’s labor supply by approximately 4.3 weeks per year. This estimate is consistent with the 2SLS benchmark.
@@ Appendix A Additional Definitions
Before the full proof, we would like to provide a high-level summary of our proof strategy to facilitate readers’ understanding. Conceptually, our proof tackles three key challenges:
  • (1)
    Two-stage coupling: We control a pair of noisy recursions $\{\bm{\Theta}^{(t)}\}_{t\leq T}$ and $\{\bm{\beta}^{(t)}\}_{t\leq T}$ where the second-stage gradient at time $t$ depends on the noisy first stage iterate $\bm{\Theta}^{(t)}$. This coupling does not appear in standard DP-SGD analyses.
  • (2)
    Noise propagation: Since we add Gaussian noise at every step and stage, the privacy accountant composes over all $2T$ iterations. To obtain a meaningful error bound for the final iterate $\bm{\beta}^{(T)}$, we need to separate the contributions of optimization, privacy, and sampling error. This decomposition gives rise to the three terms in Theorem 3.1 and explains the tradeoff pattern shown in Figure 2 and our experiments.
  • (3)
    Non-asymptotic rate: The analysis retains explicit dependence on all major parameters: sample size $n$, instrument dimension $q$, regressor dimension $p$, privacy budgets $(\rho_{1},\rho_{2})$, and confidence parameter $\tau$.
@@ F.1 Privacy for $\beta$ only
F.2 Comparison to FriendlyCore Approach
In this section, we provide a brief comparison between our two-stage gradient perturbation approach and a computationally friendly approach. Tsfadia et al. [2022] proposes a general DP meta-framework for aggregation tasks (e.g., averaging, clustering, covariance estimation) on data in a metric space. The key idea is: Given a dataset $D=\left(\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right)$ in a metric space and a ”friendship” predicate $f(\mathbf{x},\mathbf{y})$ that encodes when two points are close / well-behaved, FriendlyCore is a DP procedure that extracts a subset $C\subseteq D$ (the ”core”) with two properties:
  • Structural niceness: $C$ is $f$-friendly (e.g., all points in $C$ lie in a ball of radius $r$, or satisfy a separation condition useful for clustering). Outliers that violate the predicate are removed.
  • Stability and DP: For neighboring datasets $D,D^{\prime}$, the cores $C,C^{\prime}$ differ only on a small number of points, and FriendlyCore itself is $(\varepsilon,\delta)$-DP or $\rho$-zCDP. This lets you plug $C$ into any friendly-instance DP algorithm without re-doing a worst-case sensitivity analysis.
Section 5 of Tsfadia et al. [2022] shows how to use this framework for private averaging, clustering, and covariance estimation. For averaging, Algorithm 5.1 FC_Avg does:
  • Split the privacy budget as $\rho_{1}=0.1\rho,\rho_{2}=0.9\rho$.
  • Run FriendlyCore on $D$ with a predicate that enforces an effective diameter $r$ (all but a few outliers lie in a ball of radius $r$ ).
  • On the core $C$, run FriendlyAvg, which is essentially a Gaussian-mechanism mean estimator tuned for zCDP .
Algorithm 5.1 from Tsfadia et al. [2022] can be used as a building block to make a DP version of this 2SLS analysis: The 2SLS estimator depends on sample means of sufficient statistics: $\frac{1}{n}\mathbf{Z}^{\top}\mathbf{Z},\frac{1}{n}\mathbf{Z}^{\top}\mathbf{X},\frac{1}{n}\mathbf{Z}^{\top}\mathbf{Y}$. Stack and vectorize these matrices into a vector in $\mathbb{R}^{d}$ with $d=q^{2}+pq+q=\mathcal{O}(q^{2})$. Each data point contributes a vector of this form; call these contributions $\tilde{\mathbf{x}}_{i}$. Under the sub-Gaussian design, each $\tilde{\mathbf{x}}_{i}$ has bounded effective diameter
$r=\Theta(\sqrt{d}+\sqrt{\tau})$
with high probability. Using that $r$, Algorithm 5.1 (FC Avg) gives a $\rho$-zCDP estimate of the mean of the sufficient-statistics vector with error
$\mathcal{O}\left(\frac{r}{n}\sqrt{\frac{d}{\rho}}\right)=\mathcal{O}\left(\frac{\sqrt{d}+\sqrt{\tau}}{n}\sqrt{\frac{d}{\rho}}\right)=\mathcal{O}\left(\frac{d+\sqrt{d\tau}}{n\sqrt{\rho}}\right).$
Lemma D.7’s proof shows that $\hat{\beta}$ is a smooth function of those sample covariances. The final estimate $\hat{\bm{\beta}}_{\text{FC}}$ is then a post-processing step that applies the usual 2SLS formula to these privatized moments. If we replace the non-private moment estimates in Lemma D.7 by the FC_Avg privatized moments, the Lipschitz dependence of $\hat{\beta}$ on the moments converts the FC_Avg error into an additional term in the error bound, scaling like at the order of $L\frac{q^{2}+q\sqrt{\tau}}{n\sqrt{\rho}}$, where $L$ is the Lipschitz constant (see Lemma F.2). Hence, the error bound for a FriendlyCore-based DP 2SLS estimator would be the combination of the non-private rate from Lemma D.7 plus this additional term. In particular,
$\|\hat{\bm{\beta}}_{\text{FC}}-\bm{\beta}\|\leq\|\hat{\bm{\beta}}_{\text{FC}}-\hat{\bm{\beta}}\|+\|\hat{\bm{\beta}}-\bm{\beta}\|=\mathcal{O}\left(L\frac{q^{2}+q\sqrt{\tau}}{n\sqrt{\rho}}+\frac{\sqrt{pq}(\tau+\log(pq))}{\sqrt{n}}\right).$
Remark F.1.
Comparing this bound with our bound in Theorem 3.1, we see that the orders are similar. At the same time, the FriendlyCore-based estimator reflects a different set of algorithmic and statistical trade-offs from our DP-2S-GD method:
  • Black-box generality vs. algorithm-aware design. FriendlyCore offers a highly general subsample–aggregate framework that applies to a broad class of moment-based estimators, including 2SLS, by treating them as black-box functions of empirical moments. In contrast, DP-2S-GD directly privatizes the two-stage gradient procedure itself. This algorithm-aware approach enables us to analyze how privacy noise propagates through the optimization dynamics and to obtain an explicit privacy–iteration trade-off (e.g., the “too many iterations hurt” behavior in Figure 2), which does not emerge in one-shot subsample–aggregate pipelines.
  • Subsample-and-aggregate vs. full-data iterative updates. FriendlyCore relies on repeated subsampling to identify a stable “core” before applying the estimator on that core, which provides robustness and generality but increases computational cost and effectively reduces the sample size available to each subsample. DP-2S-GD uses all samples at every iteration with per-sample gradient clipping and a simple zCDP accountant, making it attractive in large-scale, iterative, or streaming settings where repeated subsampling may be impractical.
Lemma F.2.
Let $\mathbf{s}:=\text{vec}(\frac{\mathbf{Z}^{\top}\mathbf{Z}}{n},\frac{\mathbf{Z}^{\top}\mathbf{X}}{n},\frac{\mathbf{Z}^{\top}\mathbf{Y}}{n})$, and $\bm{\beta}(\mathbf{s})$ denotes the mapping from $\mathbf{s}$ to $\hat{\bm{\beta}}_{2SLS}$. There exists a constant $L>0$ such that for any two sets of moments $\mathbf{s}_{1},\mathbf{s}_{2}$, we have
$\|\bm{\beta}(\mathbf{s}_{1})-\bm{\beta}(\mathbf{s}_{2})\|\leq L\|\mathbf{s}_{1}-\mathbf{s}_{2}\|.$
Proof.
Let $\bm{\Sigma}_{zz}=\frac{1}{n}\mathbf{Z}^{\top}\mathbf{Z},\qquad\bm{\Sigma}_{zx}=\frac{1}{n}\mathbf{Z}^{\top}\mathbf{X},\qquad\bm{\Sigma}_{zy}=\frac{1}{n}\mathbf{Z}^{\top}\mathbf{Y}$ denote the empirical second moments and define the “moment vector”
$\mathbf{s}:=(\bm{\Sigma}_{zz},\bm{\Sigma}_{zx},\bm{\Sigma}_{zy}).$
From these moments we form the usual 2SLS normal equations
$\mathbf{G}(\mathbf{s})\bm{\beta}(\mathbf{s})=\mathbf{h}(\mathbf{s}),$
where $\mathbf{G}(\mathbf{s}):=\mathbf{X}^{\top}\mathbf{P}_{z}\mathbf{X},\quad\mathbf{h}(\mathbf{s}):=\mathbf{X}^{\top}\mathbf{P}_{z}\mathbf{Y},\quad\mathbf{P}_{z}:=\mathbf{Z}(\mathbf{Z}^{\top}\mathbf{Z})^{-1}\mathbf{Z}^{\top}.$ Hence,
$\mathbf{G}(\mathbf{s})=\bm{\Sigma}_{xz}\bm{\Sigma}_{zz}^{-1}\bm{\Sigma}_{zx},\qquad\mathbf{h}(\mathbf{s})=\bm{\Sigma}_{xz}\bm{\Sigma}_{zz}^{-1}\bm{\Sigma}_{zy},$
with $\bm{\Sigma}_{xz}=\bm{\Sigma}_{zx}^{\top}$. Thus the 2SLS estimator can be written as
$\bm{\beta}(\mathbf{s})=\mathbf{G}(\mathbf{s})^{-1}\mathbf{h}(\mathbf{s}).$
Consider two sets of moments $s_{1},s_{2}$ with corresponding
$\mathbf{G}_{i}:=\mathbf{G}(\mathbf{s}_{i}),\quad\mathbf{h}_{i}:=\mathbf{h}(\mathbf{s}_{i}),\quad\bm{\beta}_{i}:=\bm{\beta}(\mathbf{s}_{i})=\mathbf{G}_{i}^{-1}\mathbf{h}_{i}\quad(i=1,2).$
We have
$\bm{\beta}_{1}-\bm{\beta}_{2}$ $=\mathbf{G}_{1}^{-1}\mathbf{h}_{1}-\mathbf{G}_{2}^{-1}\mathbf{h}_{2}$
$=\mathbf{G}_{1}^{-1}(\mathbf{h}_{1}-\mathbf{h}_{2})+(\mathbf{G}_{1}^{-1}-\mathbf{G}_{2}^{-1})\mathbf{h}_{2}$
$=\mathbf{G}_{1}^{-1}(\mathbf{h}_{1}-\mathbf{h}_{2})+\mathbf{G}_{1}^{-1}(\mathbf{G}_{2}-\mathbf{G}_{1})\mathbf{G}_{2}^{-1}\mathbf{h}_{2},$
where in the last equality we used the identity
$\mathbf{G}_{1}^{-1}-\mathbf{G}_{2}^{-1}=\mathbf{G}_{1}^{-1}(\mathbf{G}_{2}-\mathbf{G}_{1})\mathbf{G}_{2}^{-1}.$
Taking norms and using submultiplicativity,
$\|\bm{\beta}_{1}-\bm{\beta}_{2}\|\leq\|\mathbf{G}_{1}^{-1}\|\,\|\mathbf{h}_{1}-\mathbf{h}_{2}\|+\|\mathbf{G}_{1}^{-1}\|\,\|\mathbf{G}_{2}-\mathbf{G}_{1}\|\,\|\mathbf{G}_{2}^{-1}\|\,\|\mathbf{h}_{2}\|.$ (53)
Assume (as in Assumption 2) that the population Gram matrix is well conditioned, so that on a high-probability event
$\lambda_{\min}(\mathbf{G}_{i})\geq\lambda_{0}>0\quad\Rightarrow\quad\|\mathbf{G}_{i}^{-1}\|\leq\lambda_{0}^{-1},\quad i=1,2,$
and that the moments are uniformly bounded so that $\|\mathbf{h}_{2}\|\leq C_{h}$. Here the specific formula of $\lambda_{0}$ could be referred to Lemma D.6. Moreover, $\mathbf{G}(\mathbf{s})$ and $h(\mathbf{s})$ are smooth (in fact, rational) functions of the entries of $(\bm{\Sigma}_{zz},\bm{\Sigma}_{zx},\bm{\Sigma}_{zy})$, and in a neighbourhood of the true moments there exist constants $C_{G},C_{h}>0$ such that
$\|\mathbf{G}_{2}-\mathbf{G}_{1}\|\leq C_{G}\|\mathbf{s}_{2}-\mathbf{s}_{1}\|,\qquad\|\mathbf{h}_{2}-\mathbf{h}_{1}\|\leq C_{h}\|\mathbf{s}_{2}-\mathbf{s}_{1}\|.$
Plugging these bounds into (53) yields
$\|\bm{\beta}_{1}-\bm{\beta}_{2}\|\leq\frac{1}{\lambda_{0}}C_{h}\|\mathbf{s}_{1}-\mathbf{s}_{2}\|+\frac{1}{\lambda_{0}^{2}}C_{G}C_{h}\|\mathbf{s}_{1}-\mathbf{s}_{2}\|=L\|\mathbf{s}_{1}-\mathbf{s}_{2}\|,$
where
$L:=\frac{C_{h}}{\lambda_{0}}+\frac{C_{G}C_{h}}{\lambda_{0}^{2}}.$
Hence the 2SLS map $\bm{\beta}:\mathbf{s}\mapsto\hat{\bm{\beta}}_{\text{2SLS}}(\mathbf{s})$ is Lipschitz in the sample moments:
$\|\bm{\beta}(\mathbf{s}_{1})-\bm{\beta}(\mathbf{s}_{2})\|\leq L\|\mathbf{s}_{1}-\mathbf{s}_{2}\|.$
@@ Appendix G Additional Experiments
G.1 Tuning Step Size
In this section, we empirically examine how the step sizes $\alpha$ and $\eta$ affect the convergence of Algorithm 1. Using the same setup as in Section 4.1, we fix $n=2000$ and $p=q=r=5$, and run Algorithm 1 for $T=20$ iterations, with $\rho_{1}=\rho_{2}=5$. In each plot, we vary one of $\eta,\alpha$ over its admissible range given by (3), while fixing the other step size at a sub-optimal level (close to its upper bound). The results, shown in Figure 7, indicate that as long as $\eta$ and $\alpha$ lie within the theoretically justified region, the convergence behavior is fairly insensitive to the exact step-size choice. As noted in Remark 3.2, our theoretical upper bound for $\alpha$ is slightly conservative due to the need to control the randomness introduced by the first-stage estimates $\bm{\Theta}^{(t)}$.
Figure 7: Convergence behavior under different step sizes $\eta,\alpha$. The (theoretical) upper bounds $\eta_{\text{max}}$ and $\alpha_{\text{max}}$ are given by (3). The (approximate) optimal $\eta_{\text{approx}}^{\star}$ and $\alpha_{\text{approx}}^{\star}$ are calculated according to (8). (a) Varying $\eta$ while fixing $\alpha=\alpha_{\text{max}}$. (b) Varying $\alpha$ while fixing $\eta=0.9\eta_{\text{max}}$. All the curves are averaged over 100 runs, with vertical bars representing the standard errors.
G.2 Effect of clipping threshold
In this section, we empirically examine how the clipping thresholds $\gamma_{1}$ and $\gamma_{2}$ influence the utility of Algorithm 1. Using the same setup as in Section 4.1, we fix $n=2000$ and $p=q=r=5$, and run Algorithm 1 for $T=20$ iterations under privacy budgets $\rho_{1}=\rho_{2}=5$. For simplicity, we set $\gamma_{1}=\gamma_{2}=\gamma$ and vary $\gamma$ over the range $[1,1000]$. The results are reported in Figure 8.
We observe that when $\gamma$ is set too small, the per-sample gradients are frequently clipped, causing the updates to be severely distorted and resulting in larger estimation error. As $\gamma$ increases, clipping becomes less frequent and the estimation error decreases. However, once $\gamma$ exceeds a certain level, the sensitivity of the gradients grows, which requires injecting larger noise to satisfy the target privacy budget. This increased noise leads to larger fluctuations in the final estimates. Consequently, the most effective choice of $\gamma$ is the smallest value that ensures gradient clipping does not occur with high probability.
Figure 8: Effect of clipping threshold $\gamma$ on the utility of Algorithm 1. We fix $n=2000,p=q=r=5,T=20,\rho_{1}=\rho_{2}=5$, and set $\gamma_{1}=\gamma_{2}=\gamma$. The error curve is averaged over 100 runs, with vertical bars representing the standard errors.
G.3 Convergence Rate Comparison
@@
G.4 Additional Experiments on Angrist Dataset
We provide additional experimental results on the Angrist dataset with different privacy parameters $\rho_{1},\rho_{2}$. We consider two settings of privacy parameters: (i) $\rho_{1}=0.1,\rho_{2}=0.1$; (ii) $\rho_{1}=10,\rho_{2}=10$. The results are shown in Figures 10 and 11. We observe that when $\rho_{1},\rho_{2}$ are small, the estimates of $\bm{\beta}^{(T)}$ have larger variance. When $\rho_{1},\rho_{2}$ are larger, the estimates of $\bm{\beta}^{(T)}$ are more concentrated around the expected value. In both settings, the estimates of $\bm{\beta}^{(t)}$ converge in expectation within $T=20$ iterations.
Figure 10: Results on the Angrist dataset with $T=20,\rho_{1}=0.1,\rho_{2}=0.1$. (a) Boxplot of estimated $\bm{\beta}^{(T)}$, over 1000 runs. (b) Learning paths of parameters $\bm{\beta}^{(t)},\bm{\Theta}^{(t)}$, over 1000 runs. The shaded area represents the standard error.
G.5 Experiments on Card Dataset
The Card dataset [Card, 1993] is a widely used empirical dataset in labor economics for studying the causal effect of education on earnings. In this study, the endogenous regressor is individuals’ years of schooling (educ), and the outcome variable is log earnings (lwage). There are several instruments available, most notably the college-proximity indicators (nearc2 and nearc4), which capture whether an individual grew up near a two-year or four-year college. Additional instruments include parental education—father’s and mother’s years of schooling (fatheduc and motheduc)—which provide further exogenous variation in educational attainment.
There are 2191 samples in total. We consider the following covariates: $\mathbf{Z}=$[nearc2, nearc4, fatheduc, motheduc], $\mathbf{X}=$[educ], $\mathbf{Y}=$[lwage]. We standardize each column of $\mathbf{Z}$ to have zero mean and unit variance. We run Algorithm 1 with privacy parameters $\rho_{1},\rho_{2}\in\{0.1,1,10\}$, and number of iterations $T=15$. We report the boxplot of final estimates and the learning path for $\bm{\beta}^{(t)}$. The results are shown in Figure 13. We observe that as $\rho_{1},\rho_{2}$ increase, the estimates of $\bm{\beta}^{(T)}$ become more concentrated. In all settings, the estimates of $\bm{\beta}^{(t)}$ converge in expectation within $T=10$ iterations.
[…]